# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ Acknowledgements
The authors are grateful to S. Treil for suggesting the interpretation of the $q$-function as an operator whose values are convex body valued functions for any $1<q<\infty,$ inspiring the interpretation of vector valued extensions of operators as operators whose values are convex body valued functions used in this paper. In addition, the authors are indebted to both Marcin Bownik and David Cruz-Uribe for carefully reading this article and for providing valuable feedback that improved the clarity and readability of this work.
@@ 2 Background
In this section we review the central definitions for our work. We also recall some basic facts about the various objects we deal with, in most cases without proof. References with further details are provided in this case.
2.1 Dyadic grids
For definiteness, in what follows intervals in $\mathbb{R}$ will always be assumed to be left-closed, right-open and bounded. A cube in $\mathbb{R}^{n}$ will be a set of the form $Q=I_{1}\times\ldots\times I_{n}$, where $I_{k},\leavevmode\nobreak\ k=1,\ldots,n$ are intervals in $\mathbb{R}$ of the same length, which we denote by $\ell(Q)\coloneq|I_{1}|.$ A rectangle in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ (with sides parallel to the coordinate axes) will be a set of the form $R=R_{1}\times R_{2}$, where $R_{1}$ is a cube in $\mathbb{R}^{n}$ and $R_{2}$ is a cube in $\mathbb{R}^{m}$.
A collection $\mathcal{D}$ of intervals in $\mathbb{R}$ will be said to be a dyadic grid in $\mathbb{R}$ if one can write $\mathcal{D}=\bigcup_{k\in\mathbb{Z}}\mathcal{D}_{k}$, such that the following hold:
  • for all $k\in\mathbb{Z}$, $\mathcal{D}_{k}$ forms a partition of $\mathbb{R}$, and all intervals in $\mathcal{D}_{k}$ have the same length $2^{-k}.$
  • for all $k\in\mathbb{Z}$, every $J\in\mathcal{D}_{k}$ can be written as a union of exactly $2$ intervals in $\mathcal{D}_{k+1}.$
We say that $\mathcal{D}$ is the standard dyadic grid in $\mathbb{R}$ if
$\mathcal{D}\coloneq\left\{[m2^{k},(m+1)2^{k}):\leavevmode\nobreak\ k,m\in\mathbb{Z}\right\}.$
A collection $\mathcal{D}$ of cubes in $\mathbb{R}^{n}$ will be said to be a dyadic grid in $\mathbb{R}^{n}$ if for some dyadic grids $\mathcal{D}^{1},\ldots,\mathcal{D}^{n}$ in $\mathbb{R}$ one can write $\mathcal{D}=\bigcup_{k\in\mathbb{Z}}\mathcal{D}_{k}$, where
$\mathcal{D}_{k}=\{I_{1}\times\ldots\times I_{n}:\leavevmode\nobreak\ I_{i}\in\mathcal{D}^{i},\leavevmode\nobreak\ |I_{i}|=2^{-k},\leavevmode\nobreak\ i=1,\ldots,n\}.$
We say that $\mathcal{D}$ is the standard dyadic grid in $\mathbb{R}^{n}$ if
$\mathcal{D}\coloneq\{[m_{1}2^{k},(m_{1}+1)2^{k})\times\dots\times[m_{n}2^{k},(m_{n}+1)2^{k}):\leavevmode\nobreak\ k,m_{1},\ldots,m_{n}\in\mathbb{Z}\}.$
If $\mathcal{D}$ is a dyadic grid in $\mathbb{R}^{n}$, then we denote
$\mathrm{ch}_{i}(Q)\coloneq\{K\in\mathcal{D}:\leavevmode\nobreak\ K\subseteq Q,\leavevmode\nobreak\ |K|=2^{-in}|Q|\},\qquad Q\in\mathcal{D},\leavevmode\nobreak\ i=0,1,2,\ldots.$
A collection $\bm{\mathcal{D}}$ is said to be a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ if for a dyadic grid $\mathcal{D}^{1}$ in $\mathbb{R}^{n}$ and a dyadic grid $\mathcal{D}^{2}$ in $\mathbb{R}^{m}$ we have
$\bm{\mathcal{D}}\coloneq\{R_{1}\times R_{2}:\leavevmode\nobreak\ R_{i}\in\mathcal{D}_{i},\leavevmode\nobreak\ i=1,2\},$
and in this case we write (slightly abusing the notation) $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$.
If $\bm{\mathcal{D}}$ is a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, then we denote
$\mathrm{ch}_{i}(R)\coloneq\{Q_{1}\times Q_{2}:\leavevmode\nobreak\ Q_{i}\in\mathrm{ch}_{i_{j}}(R_{j}),\leavevmode\nobreak\ j=1,2\},$
where
$R\in\bm{\mathcal{D}},\leavevmode\nobreak\ i=(i_{1},i_{2}),\leavevmode\nobreak\ i_{1},i_{2}=0,1,2,\ldots.$
We say that $R$ is the $i$-th ancestor of $P$ in $\bm{\mathcal{D}}$ if $P\in\mathrm{ch}_{i}(R)$.
2.2 Haar systems
2.2.1 Haar system on $\mathbb{R}$
@@
Let $\mathcal{D}$ be a dyadic grid in $\mathbb{R}$. For any interval $I\in\mathcal{D}$ we denote by $h^{0}_{I},h^{1}_{I}$ the $\mathrm{L}^{2}$-normalized cancellative and noncancellative respectively Haar functions over the interval $I\in\mathcal{D}$, that is
$h^{0}_{I}\coloneq\frac{\mathbf{1}_{I_{+}}-\mathbf{1}_{I_{-}}}{\sqrt{|I|}},\qquad h^{1}_{I}\coloneq\frac{\mathbf{1}_{I}}{\sqrt{|I|}}$
(so $h^{0}_{I}$ has mean 0). For simplicity we denote $h_{I}\coloneq h_{I}^{0}$. For any function $f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R})$, we denote $f_{I}\coloneq(f,h_{I})$, $I\in\mathcal{D}$. It is well-known that the system $\{h_{I}\}_{I\in\mathcal{D}}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R})$. Of course, all these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions in the obvious entrywise way.
2.2.2 Haar system on $\mathbb{R}^{n}$
Let $\mathcal{D}$ be a dyadic grid in $\mathbb{R}^{n}$. We denote
$\mathcal{E}\coloneq\{0,1\}^{n}\setminus\{(1,\ldots,1)\}.$
We call $\mathcal{E}$ the set of one-parameter signatures. For a cube $Q=I_{1}\times\dots\times I_{n}\in\mathcal{D}$ and $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{n})\in\mathcal{E},$ we denote by $h^{\varepsilon}_{Q}$ the $\mathrm{L}^{2}$-normalized cancellative Haar function over the cube $Q$ corresponding to the signature $\varepsilon$ defined by
$h^{\varepsilon}_{Q}(x)\coloneq h^{\varepsilon_{1}}_{I_{1}}(x_{1})\dots h^{\varepsilon_{n}}_{I_{n}}(x_{n}),\qquad x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}.$
For any function $f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n})$ we denote $f^{\varepsilon}_{Q}\coloneq(f,h^{\varepsilon}_{Q})$. It is well-known that the system $\{h^{\varepsilon}_{Q}\colon\leavevmode\nobreak\ Q\in\mathcal{D},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}\}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R}^{n})$. All these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions in the obvious entrywise way.
2.2.3 Haar system on the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ We denote $\mathcal{E}\coloneq\mathcal{E}^{1}\times\mathcal{E}^{2}$, where
$\mathcal{E}^{1}\coloneq\{0,1\}^{n}\setminus\{(1,\ldots,1)\},\leavevmode\nobreak\ \mathcal{E}^{2}\coloneq\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$
We call $\mathcal{E}$ the set of biparameter signatures. For $R=R_{1}\times R_{2}\in\bm{\mathcal{D}}$ and $\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\mathcal{E},$ we denote by $h_{R}^{\varepsilon}$ the $\mathrm{L}^{2}$-normalized cancellative Haar function over the rectangle $R$ corresponding to the signature $\varepsilon$ defined by
$h_{R}^{\varepsilon}=h_{R_{1}}^{\varepsilon_{1}}\otimes h_{R_{2}}^{\varepsilon_{2}},$
that is
$h_{R}^{\varepsilon}(x_{1},x_{2})=h_{I}^{\varepsilon_{1}}(x_{1})h_{J}^{\varepsilon_{2}}(x_{2}),\qquad(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$
For any function $f\in\mathrm{L}_{\mathrm{loc}}^{1}(\mathbb{R}^{n+m})$ we denote $f_{R}^{\varepsilon}\coloneq(f,h_{R}^{\varepsilon})$. From the corresponding one-parameter facts, we immediately deduce that the system $\{h_{R}^{\varepsilon}\colon R\in\bm{\mathcal{D}},\varepsilon\in\mathcal{E}\}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R}^{n+m}).$ For $P\in\mathcal{D}^{1},$ $Q\in\mathcal{D}^{2}$, $\varepsilon_{1}\in\mathcal{E}^{1}$ and $\varepsilon_{2}\in\mathcal{E}^{2}$ we denote
$f^{\varepsilon_{1},1}_{P}(x_{2})\coloneq(f(\,\cdot\,,x_{2}),h_{P}^{\varepsilon_{1}}),\quad x_{2}\in\mathbb{R}^{m},$
$f^{\varepsilon_{2},2}_{Q}(x_{1})\coloneq(f(x_{1},\,\cdot\,),h_{Q}^{\varepsilon_{2}}),\quad x_{1}\in\mathbb{R}^{n}.$
All these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions entrywise.
Finally, we remark that the dimensions $n,m$ will always be clear from the context, as well as whether $\mathcal{E}$ refers to the set of one-parameter or biparameter signatures.
2.3 Matrix norms in terms of column norms
In the sequel we denote by $\{e_{1},\ldots,e_{d}\}$ the standard basis of $\mathbb{C}^{d}.$ We will be often using the fact that
$|A|\sim_{d}\sum_{k=1}^{d}|Ae_{k}|,\quad\forall A\in\mathrm{M}_{d}(\mathbb{C}),$ (2.1)
without explicitly mentioning it. Note also that for all $0<p<\infty$ and for nonnegative numbers $x_{1},\ldots,x_{d}$ we have the estimate
$\min(1,d^{p-1})\sum_{i=1}^{d}x_{i}^{p}\leq\left(\sum_{i=1}^{d}x_{i}\right)^{p}\leq\max(1,d^{p-1})\sum_{i=1}^{d}x_{i}^{p}.$
In particular, if $F_{k}:\mathbb{R}^{n}\rightarrow\mathrm{M}_{d}(\mathbb{C})$ is a sequence of Lebesgue-measurable functions and $0<p,q,r<\infty,$ then we have
$\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}|F_{k}(x)|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q,r}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}\left(\sum_{i=1}^{d}|F_{k}(x)e_{i}|\right)^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q,r}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}\sum_{i=1}^{d}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q}\left(\int_{\mathbb{R}^{n}}\sum_{i=1}^{d}\left(\sum_{k=1}^{\infty}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,q}\sum_{i=1}^{d}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}.$ (2.2)
2.4 Matrix weighted matrix-valued Lebesgue spaces
A function $W$ on $\mathbb{R}^{n}$ is said to be a $d\times d$-matrix valued weight, or just matrix weight, if it is a locally integrable $\mathrm{M}_{d}(\mathbb{C})$-valued function such that $W(x)$ is a positive-definite matrix for a.e. $x\in\mathbb{R}^{n}.$ Here, by locally integrable matrix valued function $W(x)$ we mean that the scalar valued function $|W(x)|,$ that is the matrix norm, is locally integrable.
Given a $d\times d$-matrix weight $W$ on $\mathbb{R}^{n}$ and $1<p<\infty,$ we define the norm
$\|F\|_{\mathrm{L}^{p}(W)}\coloneq\left(\int_{\mathbb{R}^{n}}|W(x)^{1/p}F(x)|^{p}\,\mathrm{dm}(x)\right)^{1/p},$
for all $\mathrm{M}_{d}(\mathbb{C})$-valued measurable functions $F$ on $\mathbb{R}^{n}.$ This norm defines the matrix weighted matrix-valued Lebesgue space $\mathrm{L}^{p}(W).$
@@ 2.5 Reducing operators
Let $1<p<\infty$. Let $E$ be a bounded measurable subset of $\mathbb{R}^{n}$ of nonzero measure. Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ that is integrable over $E$ (meaning that its matrix norm is integrable, so $\int_{E}|W(x)|\,\mathrm{dm}(x)<\infty$) and that takes a.e. values in the set of positive-definite $(d\times d)$ matrices. It is proved in [Goldberg_2003, Proposition 1.2] that there exists a (not necessarily unique) positive-definite matrix $\mathcal{W}_{E}\in\mathrm{M}_{d}(\mathbb{C})$, called reducing operator of $W$ over $E$ with respect to the exponent $p$, such that
$\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}e|^{p}\,\mathrm{dm}(x)\right)^{1/p}\leq|\mathcal{W}_{E}\,e|\leq\sqrt{d}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}e|^{p}\,\mathrm{dm}(x)\right)^{1/p}$
for all $e\in\mathbb{C}^{d}.$ If $d=1$, i.e. $W$ is scalar-valued, then one can clearly take $\mathcal{W}_{E}\coloneq\langle W\rangle_{E}^{1/p},$ where we denote the average of $W$ over $E$ by
$\langle W\rangle_{E}=\frac{1}{|E|}\int_{E}W(x)\,\mathrm{dm}(x).$
Moreover, if $p=2$, then
$\frac{1}{|E|}\int_{E}|W(x)^{1/2}e|^{2}\,\mathrm{dm}(x)=\frac{1}{|E|}\int_{E}\langle W(x)e,e\rangle\,\mathrm{dm}(x),$
which by linearity is equal to
$\langle\langle W\rangle_{E}\,e,e\rangle=|\langle W\rangle_{E}^{1/2}e|^{2}$
for all $e\in\mathbb{C}^{d}.$ Thus in this special case one can take $\mathcal{W}_{E}\coloneq\langle W\rangle_{E}^{1/2}.$
@@
Assume in addition now that the function $W^{\prime}\coloneq W^{-1/(p-1)}$ is also integrable over $E$. Then we let $\mathcal{W}^{\prime}_{E}$ be the reducing matrix of $W^{\prime}$ over $E$ corresponding to the exponent $p^{\prime}\coloneq p/(p-1)$, so that
$|\mathcal{W}^{\prime}_{E}\,e|\sim_{d}\left(\frac{1}{|E|}\int_{E}|W^{\prime}(x)^{1/p^{\prime}}e|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p^{\prime}}=\left(\frac{1}{|E|}\int_{E}|W(x)^{-1/p}e|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p^{\prime}}$
for all $e\in\mathbb{C}^{d}.$ Note that one can take $\mathcal{W}^{\prime\prime}_{E}=\mathcal{W}_{E}$. Observe that
$|\mathcal{W}_{E}\mathcal{W}^{\prime}_{E}|\sim_{p,d}\left(\frac{1}{|E|}\int_{E}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p}.$
For a detailed exposition of reducing operators we refer for example to [DKPS2024]. Here we just state the following estimates for later convenience.
Lemma 3.
Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ taking a.e. positive-definite values, and such that $W$ and $W^{\prime}=W^{-1/(p-1)}$ are integrable over $E$ for some $1<p<\infty.$ Set
$C_{E}\coloneq\frac{1}{|E|}\int_{E}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x).$
We consider reducing operators of $W$ with respect to exponent $p$ and of $W^{\prime}$ with respect to exponent $p^{\prime}.$
  1. (1)
    There holds
    $|\mathcal{W}_{E}e|\lesssim_{p,d}C_{E}^{1/p}{\langle|W^{1/p}e|\rangle}_{E}$
    and
    $|\mathcal{W}_{E}^{-1}e|\leq|\mathcal{W}_{E}^{\prime}e|\lesssim_{p,d}C_{E}^{1/p}|\mathcal{W}_{E}^{-1}e|,$
    for all $e\in\mathbb{C}^{d}.$
  2. (2)
    There holds
    $|{\langle W^{1/p}\rangle}_{E}e|\leq|\mathcal{W}_{E}e|\lesssim_{p,d}C_{E}^{d/p}|{\langle W^{1/p}\rangle}_{E}e|,$
    for all $e\in\mathbb{C}^{d}.$
A proof of part (1) can be found, for example, in [DKPS2024]. A proof of part (2) can be found in [Isralowitz_Kwon_Pott2017].
@@
Let $E,F$ be measurable subsets of $\mathbb{R}^{n},\mathbb{R}^{m}$ respectively with $0<|E|,|F|<\infty.$ Let $1<p<\infty.$ Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued integrable function on $E\times F$ taking a.e. values in the set of positive-definite $d\times d$-matrices. For all $x_{1}\in E,$ set $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in F.$ For a.e. $x_{1}\in E,$ denote by $\mathcal{W}_{x_{1},F}$ the reducing operator of $W_{x_{1}}$ over $F$ with respect to the exponent $p.$ It is proved in [DKPS2024] (see also [Cruz_Uribe_Bownik_Extrapolation]) that one can choose the reducing operator $\mathcal{W}_{x_{1},F}$ in a way that is measurable in $x_{1}.$
Set $W_{F}(x_{1})\coloneq\mathcal{W}_{x_{1},F}^{p},$ for a.e. $x_{1}\in E.$ Then $W_{F}\in\mathrm{L}^{1}(E;\mathrm{M}_{d}(\mathbb{C})).$ It is proved in [DKPS2024] that
@@
where $\mathcal{W}_{F,E}$ is the reducing operator of $W_{F}$ over $E$ with respect to the exponent $p,$ and $\mathcal{W}_{E\times F}$ is the reducing operator of $W$ over $E\times F$ with respect to the exponent $p.$

### v3

@@ Acknowledgements
The authors are grateful to S. Treil for suggesting the interpretation of the $q$-function as an operator whose values are convex body valued functions for any $1<q<\infty,$ inspiring the interpretation of vector valued extensions of operators as operators whose values are convex body valued functions used in this paper. In addition, the authors are indebted to both Marcin Bownik and David Cruz-Uribe for carefully reading this article and for providing valuable feedback that improved the clarity and readability of this work. Finally, the authors are also thankful to the anonymous referee, who made very useful suggestions and comments to improve the clarity of this article.
@@ 2 Background
2.1 Notation for dyadic grids and Haar systems
We briefly review the notation that we use for product dyadic grids. Consider two different dyadic grids $\mathcal{D}^{1}$ and $\mathcal{D}^{2}$ on $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$ respectively; that is, both $\mathcal{D}^{1}$ and $\mathcal{D}^{2}$ are collections of dyadic cubes. Then we say that $\bm{\mathcal{D}}$ is the product dyadic grid given by $\mathcal{D}^{1}$ and $\mathcal{D}^{2}$ on $\mathbb{R}^{n}\times\mathbb{R}^{m},$ denoted by $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2},$ if $\bm{\mathcal{D}}$ is the collection of dyadic rectangles of the form $R_{1}\times R_{2}$ with $R_{j}\in\mathcal{D}^{j},$ $j=1,2.$
@@ 2.1 Notation for dyadic grids and Haar systems
Given a dyadic interval $I$ on $\mathbb{R},$ we denote the cancellative and noncancelative Haar functions on $I$ by
$h^{0}_{I}\coloneq\frac{\mathbf{1}_{I_{+}}-\mathbf{1}_{I_{-}}}{\sqrt{|I|}},\qquad h^{1}_{I}\coloneq\frac{\mathbf{1}_{I}}{\sqrt{|I|}}.$
Consider a dyadic grid $\mathcal{D}$ on $\mathbb{R}^{n}$ and let $\mathcal{E}=\{0,1\}^{n}\setminus\{(1,\ldots,1)\}.$ The Haar basis on $\mathbb{R}^{n}$ adapted to $\mathcal{D}$ is the set of functions indexed by $Q\in\mathcal{D}$ and $\varepsilon\in\mathcal{E}$ given by
$h^{\varepsilon}_{Q}(x)\coloneq h^{\varepsilon_{1}}_{I_{1}}(x_{1})\dots h^{\varepsilon_{n}}_{I_{n}}(x_{n}),$
where $Q=I_{1}\times\dots\times I_{n}$ ($\mathcal{E}$ is called the signature set of signatures on $\mathbb{R}^{n}$). Similarly, in the biparameter context, consider a product dyadic grid $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ and the set of biparameter signatures $\mathcal{E}=\mathcal{E}^{1}\times\mathcal{E}^{2},$ where $\mathcal{E}^{1}$ and $\mathcal{E}^{2}$ are respectively the signature sets on $\mathbb{R}^{n}$ and $\mathbb{R}^{m}.$ The biparameter Haar basis on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ adapted to $\bm{\mathcal{D}}$ is the set of functions
$h^{\varepsilon}_{R}(x)\coloneq h^{\varepsilon_{1}}_{R_{1}}(x_{1})\cdot h^{\varepsilon_{2}}_{R_{2}}(x_{2}),$
with $R=R_{1}\times R_{2}\in\bm{\mathcal{D}}$ and $\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\mathcal{E}.$
For a given locally integrable function $f,$ we will denote the Haar coefficient of $f$ corresponding to $h^{\varepsilon}_{R}$ by $f^{\varepsilon}_{R}$ (with analogous notation for the one parameter setting). In other words, it is the coefficient defined by
$f^{\varepsilon}_{R}=(f,h^{\varepsilon}_{R}).$ (2.1)
We will denote the average of $f$ on a rectangle $R$ by $\langle f\rangle_{R}.$
@@ 2.2 Reducing operators
Let $1<p<\infty$. Let $E$ be a bounded measurable subset of $\mathbb{R}^{n}$ of nonzero measure. Let $W$ be an integrable $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ taking a.e. positive-definite values. Here, by a positive-definite matrix $M\in\mathrm{M}_{d}(\mathbb{C})$ we mean a matrix such that $\langle Me,e\rangle\geq 0$ for every $e\in\mathbb{C}^{d}.$ In particular, since we are focusing on the space $\mathbb{C}^{d}$, note that this implies that a positive-definite matrix $M$ is also Hermitian. It is proved in [20, Proposition 1.2] that there exists a (not necessarily unique) positive-definite matrix $\mathcal{W}_{E}\in\mathrm{M}_{d}(\mathbb{C})$, called reducing operator of $W$ over $E$ with respect to the exponent $p$, such that
$\langle|W^{1/p}e|\rangle_{E,p}\leq|\mathcal{W}_{E}\,e|\leq\sqrt{d}\langle|W^{1/p}e|\rangle_{E,p}$
@@
for all $e\in\mathbb{C}^{d},$ where $\langle f\rangle_{E,p}$ denotes the $\mathrm{L}^{p}$ average of a scalar valued function $f$ over $E$. If the function $W^{\prime}\coloneq W^{-1/(p-1)}$ is also integrable over $E,$ then we let $\mathcal{W}^{\prime}_{E}$ be the reducing matrix of $W^{\prime}$ over $E$ corresponding to the exponent $p^{\prime}\coloneq p/(p-1)$. For a detailed exposition of reducing operators we refer for example to [14].
@@
Let $E,F$ be bounded measurable subsets of $\mathbb{R}^{n},\mathbb{R}^{m}$ respectively of nonzero measure. Let $1<p<\infty.$ Let $W$ be an integrable $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E\times F$ taking a.e. positive definite values. For a.e. $x_{1}\in E,$ denote by $\mathcal{W}_{x_{1},F}$ the reducing operator of $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2})$ over $F$ with respect to the exponent $p.$ It is proved in [14] (see also [5]) that one can choose the reducing operator $\mathcal{W}_{x_{1},F}$ in a way that is measurable in $x_{1}.$ Moreover, it is shown in [14] that
@@
where $\mathcal{W}_{F,E}$ is the reducing operator of $W_{F}(x_{1})\coloneq\mathcal{W}_{x_{1},F}^{p}$ over $E$ with respect to the exponent $p,$ and $\mathcal{W}_{E\times F}$ is the reducing operator of $W$ over $E\times F$ with respect to the exponent $p.$